OpenStudy (campayne):

please help someone 1.) For the following function, evaluate f(0), f(1.5), and f(-4). f(x) = 3x² - x + 2 2.) Find f(3) of f(x) = -5x + 9. Enter a numerical answer; example: -10.

OpenStudy (kailey7277):

1.) \[f(0)=3(0)^{2}-0+2\] \[f(0)=0^{2}-0+2\] \[f(0)=0-0+2\] \[f(0)=2\] \[f(1.5)=3(1.5)^{2}-1.5+2\] \[f(1.5)=4.5^{2}-1.5+2\] \[f(1.5)=20.25-1.5+2\] \[f(1.5)=20.75\] \[f(-4)=3(-4)^{2}-(-4)+2\] \[f(-4)=-12^{2}+4+2\] \[f(-4)=144+6\] \[f(-4)=150\]

OpenStudy (kailey7277):

2.) \[f(3)=-5(3)+9\] \[f(3)=-15+9\] \[f(3)=-6\] \[Answer: -6\]

OpenStudy (campayne):

you're my hero

OpenStudy (kailey7277):

haha sweet

OpenStudy (campayne):

wait but these are my options so idk -2, 3.25, 50 2, 7.25, 54 0, 5.25, 52 0, 3.25, 54

OpenStudy (campayne):

@kailey7277

OpenStudy (kailey7277):

I guess I messed up loll

OpenStudy (kailey7277):

idk im pretty sure I did it right

OpenStudy (jackthegreatest):

@CamPayne @kailey7277 kailey did both 2 and 3 wrong, remember the order of math, exponents first

OpenStudy (jackthegreatest):

@kailey7277 ...you might want to correct 2 and 3